The equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the U.S. x years after 2000,that is,x=0 corresonds to 2000,x=2 corresonds to 2002,x=6 corresonds to 2006, and so on. Predict the number of gun deaths in 2006 and 2009. In wagt year will the number of gun deaths be 9383?
The predicted gun deaths in 2006 will be?
the predicted gun deaths in 20069 will be?
The predicted number of gun deaths will be 9383 in the year?

Found 3 solutions by jorel1380, stanbon, Alan3354:
Answer by jorel1380(3719)

y=-1777(6)+27153
y=16,491
------------
y=-1777(9)+27153
y=11160
-------------------------
9383=-1777x+27153
1777x=17770
x=10
Gun deaths will be 9383 in 2010..

Answer by stanbon(75887)


The function H described by H(x)=2.75x+71.48 can be used to predict the height,in centimeters,of a woman whose humerus is x cm long.
Predict the height of a woman whose humerus is 38 cm long.
-----------
H(38) = 2.75*38+71.48 = 175.98 cm
=====================================
cheers,
Stan H.

Answer by Alan3354(69443)

Making up a formula doesn't mean anyone can predict the future.
---------------
•Homicide: 11,624 / 39% of All Fatalities
•Suicide: 16,750 / 57% of All Fatalities
•Unintentional Death (Accidental): 649 / 2% of All Fatalities
-------------
A total of 28843
-----------------------
A sad, but interesting, fact is that homocide is the #1 cause of death of pregnant women in the US (not necessarily from guns).